# Lattice-Renormalized Tunneling Models for Superconducting Qubit Materials

### v1

@@ Introduction.
The aforementioned limitations are well documented. Ring-polymer calculations show that the MEP is not necessarily the most efficient tunneling path, implying that tunnel splittings derived from it systematically underestimate transition frequencies [17]. Moreover, a 1D Schrödinger equation with a fixed effective mass is ill-defined along a path with continuously varying nuclear coordinates. Similarly, the light-particle approach assumes symmetrized structures that are energetically unrealistic: for example, the symmetrized structures used to model hydrogen TLS in bcc Nb exhibit formation energies exceeding 10 meV (Table 1, $E_{f}=E_{c}>0$); values far above the thermal energy available in quantum circuits operating below 100 mK. Even larger formation energies (>30>30 meV) have been reported for symmetrized structures in Al2O3 [16]. These shortcomings motivate the need for a tunneling model that rigorously incorporates lattice degrees of freedom while preserving thermodynamic stability.
@@
In this Letter, we introduce a formalism derived directly from the nuclear Hamiltonian that overcomes limitations of existing minimum-energy-path and light-particle approaches. Our method incorporates composite phonon coordinates to capture lattice distortions between degenerate wells, enabling a rigorous treatment of tunneling dynamics. Applied to hydrogen TLS in bcc Nb, the formalism produces tunnel splittings that bound experimental transition frequencies from below and above and reveals a fundamental connection between configurational TLS and lattice phonons beyond conventional strain coupling. It further generalizes to H multi-level systems arising from clusters of three or more spatially-localized, degenerate, configurations. We conclude by discussing the implications of our results for defect mitigation in superconducting qubit applications.
@@
Defect Structures.
A prior work has demonstrated that interstitial, self-trapped H favors tetrahedral sites in bcc Nb [18]; however, at low temperatures, self-trapped H precipitates into ordered niobium hydrides [19, 20, 21, 9]. Consequently, most H atoms that exhibit tunneling behavior must be trapped by other point defects such as O, Ti, and Zr [22, 23, 24, 25]. We apply our formalism to self-trapped H and H trapped by adjacent O, Ti, and Zr. Oxygen, which occupies interstitial octahedral sites in the bcc structure [26], is well-known to exhibit hydrogen tunneling behavior [23, 24], whereas Ti and Zr are substitutional defects that also exhibit hydrogen tunneling behavior [22, 27, 25].
Theory.
@@
Here, we ignore the cross-coupling to the remaining phonon modes $\{\omega_{iQk}\}_{i\neq\mathbf{q},Q}$ as they weakly influence the computed tunnel splitting, as we argue below. This formulation retains the essential anharmonicity required for accurate tunneling predictions and avoids reliance on thermodynamically unstable symmetrized structures. Physically, the dominant anharmonicity arises from the hydrogen phonon modes and the composite coordinate $Q$, which parameterizes the Nb lattice distortion between degenerate configurations (Fig. 1, left) and discussed further below. The full derivation, including the transformation from the nuclear Hamiltonian and its expansion in phonon coordinates, is provided in the Appendix. Note that in the limit where $Q$ is fixed at its symmetric value, $Q_{s}$, Eq. 1 reduces to a 3D H coordinate spanping both wells. (The conventional light-particle picture in a rigid lattice.)
@@
1/16111Computed using 5D Hamiltonian. 1.32 0.069 18.4 20.7222Evaluated using grid interpolation. 275.411footnotemark: 1 245.3 - $1.9\times 10^{-3}$ -
@@
Numerical Implementation.
[…]
Localization and Lattice Coupling.
@@
The wavefunction of the first-excited state follows a similar trend; however, it undergoes a phase flip at the symmetric configuration that persists toward the right well. Fig. 1g (h) further shows that the potential energy and wavefunction along the $Q$ coordinate has a harmonic character at each H minima, $\mathbf{q}_{1}$ ($\mathbf{q_{2}}$). We also plot the image of the potential energy surface at the other H minima, $\mathbf{q}_{2}$ ($\mathbf{q}_{1}$). Both surfaces meet at common coordinate, $Q_{c}$, and potential energy, $E_{c}^{\prime}$, which defines the symmetric configuration and effective coincidence energy.
Tunnel Splitting and Dimensionality Effects.
With the full 4D Hamiltonian, we compute a tunnel splitting ($J_{\mathrm{H,D}}$) ranging from 0.31 meV ($c=1/16$) to 0.064 meV ($c=1/54$) for hydrogen (H) and 0.024 meV ($c=1/16$) to $4.8\times 10^{-3}$ meV ($c=1/54$) for deuterium (D) TLS in bcc Nb (Table 1). $J_{\mathrm{H}}$ computed at $c=1/54$ is approximately 3X smaller than experimental estimates derived from specific heat (0.19 meV) and neutron scattering experiments (0.21 meV) [23], consistent with later measurements reporting values ranging from 0.170 – 0.230 meV [24]. $J_{\mathrm{D}}$ computed at $c=1/54$ is about 4 times lower than the estimate from specific heat experiments (0.021 meV) [29]. Interestingly, $J_{H}$ for self-trapped hydrogen is comparable to $J_{D}$ computed for oxygen-trapped hydrogen. This arises due to the confluence of a larger H–H site separation ($R_{H}$), a larger lattice displacement ($R_{c}$) and coincidence energy ($E_{c}$) between the local and symmetric configurations, and a larger barrier ($V$) in the symmetric configuration.
@@
Given the sensitivity of the tunnel splitting to the anharmonic terms in the Hamiltonian, we have further validated our results by comparing the computed transition frequencies between the ground state and excited states against measured inelastic neutron scattering data for O-H and H defects [19]. The reported values are 108 (106) meV and 159 (163) meV for O-H defect pairs and self-trapped H, with the latter transition two-fold degenerate due to the local symmetry of tetrahedral sites. Our results reproduce these values within a few meV at both defect concentrations (Table 3). By including the composite phonon mode $Q$, we find additional transitions appear in the excitation spectrum of the 4D Hamiltonian with frequencies dictated by the nearly harmonic states of the $Q$ mode  [^4] As the energy of these excited states overlap with the excited-state energies of the H-dominant modes, assessing the appropriate transition frequencies required a manually assignment. This was performed by plotting the wavefunctions of the excited states and selecting the excited states whose nominal wavefunction was consistent with the harmonic oscillator wavefunction of the desired transition.. The average transition energy between the ground and first excited state varied between 103-110 meV for all evaluated defects, indicating a weak sensitivity to barrier height. The average transition between the ground state and higher quasi-degenerate excited states ranged between 140–173 meV depending on $c$ and defect type. These results confirm that our sampled potential $V(\mathbf{q},Q)$ accurately captures the relevant O-H energy landscape and represent a clear improvement over the prior 3D Hamiltonian approach [18], likely due to the higher fidelity DFT functional choice and denser sampling.
@@
Impact of Nb Mode Coupling on Tunnel Splitting.
The effect of the harmonic Nb mode couplings on $J$ is expected to be small. At the symmetric configuration with a potential energy minima at $Q=Q_{s}$, the wells are degenerate, however, second-order mode couplings $\omega_{q_{i}q_{Q}}$ to the phonon bath slightly offset and lower the local minimum from $\{q_{i}\}_{\neq Q}=0$. Degenerate wells remain following this offset, which reduces the transition energy and marginally increases the tunnel splitting. We estimate the maximum reduction by comparing the energy difference between the effective, $E_{c}^{\prime}$, and true, $E_{c}$, coincidence energies when all other modes are fully relaxed at $Q=Q_{s}$. This energy difference is typically between 121-2 meV lower than the symmetric structure, corresponding to $\sim 10-20\%$ of the total symmetrization energy. For the O-H defect, $E_{c}$ is only 1 meV lower than $E_{c}^{\prime}$; such a minor correction cannot reconcile our computed $J=0.064$ meV ($c=1/54$) with the experimental value of $J_{\mathrm{H}}=0.2$ meV. Indeed, this result follows from the fact that $E_{c}^{\prime}$ for $c=1/16$ is 5 meV lower than $E_{c}^{\prime}$ for $c=1/54$, yet $J_{\mathrm{H}}$ is only 50% larger than the experimental value.
@@
Mass Effects.
The dominant anharmonicity arises from the cross-coupling $\mathbf{q}-Q$ between the H modes and the composite Nb lattice mode. This interaction strongly influences tunneling behavior and challenges the validity of the light-particle picture. For a harmonic oscillator, the wavefunction amplitude $\psi\propto\exp{(-\omega Q^{2}/(2\hbar))}$. The mode frequency along the composite coordinate is given by $\omega=\sqrt{2E_{c}}/Q_{c}$, which leads to the ratio: $\psi(Q_{c})/\psi(0)=\exp\left(-\sqrt{2E_{c}}/(2\hbar\omega)\right)$. By scaling $Q$ over a wide range of values, we evaluate how the mode frequency and coincidence energy affect the tunneling frequency of O-H and O-D defects. This is readily achieved by replacing $Q$ with $\sqrt{m/m_{\mathrm{Nb}}}Q$. Physically, this rescaling either enhances ($m<m_{\mathrm{Nb}}$) or suppresses ($m>m_{\mathrm{Nb}}$) tunneling by increasing (decreasing) the magnitude of the wavefunction in the symmetric configuration ($Q=Q_{s}$) and thereby increasing (decreasing) wavefunction overlap. Fig. 2 shows that tunnel splitting decreases by approximately a factor of 2-3 when moving from Nb to Ta. We also find that $J=J_{\mathrm{4D}}$ can vary by nearly an order of magnitude compared to the light-particle value ($J_{\mathrm{3D}})$, highlighting the critical role of these anharmonic couplings in renormalizing tunnel splittings. This reduction shifts the expected resonant frequency of H-based TLS into the 1–20 GHz range. However, strain coupling energies of 0.2–1 eV imply that even small lattice strains (as low 0.002-0.02%) are sufficient to raise the resonant frequency above typical superconducting qubit operating frequencies, effectively suppressing most TLS from acting as resonant defects.
@@
Acoustic Phonons and Strain Interactions.

### v2

@@ I Introduction
The aforementioned limitations are well documented. Ring-polymer calculations show that the MEP is not necessarily the most efficient tunneling path, implying that tunnel splittings derived from it should systematically underestimate transition frequencies [17]. Moreover, a 1D Schrödinger equation with a fixed effective mass is ill-defined along a path with continuously varying nuclear coordinates. Similarly, the light-particle approach assumes symmetrized structures that are energetically unrealistic: for example, the symmetrized structures used to model hydrogen TLS in bcc Nb exhibit formation energies exceeding 10 meV ($E_{f}=E_{c}>0$, see below); values far above thermal energy available in quantum circuits operating below 100 mK. Even larger formation energies (>30>\!30 meV) were reported for symmetrized structures in Al2O3 [16]. These shortcomings motivate the need for a tunneling model that rigorously incorporates lattice degrees of freedom while preserving thermodynamic stability.
@@
Here we introduce a formalism derived directly from the nuclear Hamiltonian that overcomes limitations of existing minimum-energy-path and light-particle approaches. Our method incorporates composite phonon coordinates to capture lattice distortions between degenerate wells, enabling a rigorous treatment of tunneling dynamics. Applied to hydrogen TLS in bcc Nb, the formalism produces tunnel splittings that bound experimental transition frequencies from below and above and reveals a fundamental connection between configurational TLS and lattice phonons beyond conventional strain coupling. It further generalizes to H multi-level systems arising from clusters of three or more spatially-localized, degenerate, configurations. We conclude by discussing the implications of our results for defect mitigation in superconducting qubit applications.
Figure 1: (a, b) H interstitial sites in BCC Nb within $2\times 2\times 1/2$ supercell. The out-of-plane position is indicated by solid ($z=0$), dash-dotted ($z=a/4$), and dashed lines ($a/2$). H sites are labeled and shaded by their distance from an (a) interstitial O site or a (b) substitutional Ti site.
II Materials System and Defect Sites
A prior work has demonstrated that interstitial, self-trapped H favors tetrahedral sites in bcc Nb [18]; however, at low temperatures, self-trapped H precipitates into ordered niobium hydrides [19, 20, 21, 9]. Consequently, most H atoms that exhibit tunneling behavior must be trapped by other point defects such as O, Ti, and Zr [22, 23, 24, 25]. We therefore apply our formalism to self-trapped H and H trapped by adjacent O, Ti, and Zr. Oxygen, which occupies interstitial octahedral sites in the bcc structure [26], is well-known to exhibit hydrogen tunneling behavior [23, 24], whereas Ti and Zr are substitutional defects that also exhibit hydrogen tunneling behavior [22, 27, 25]. We also consider the potential for H to be trapped by substitutional Ta whose presence in Nb films may be expected due to its chemical similarity to Nb and possible role as a encapsulation layer for Nb films [6].
Table 1: Formation energy ($E_{f}$) of an O-H complex in bcc Nb with H located at the 1st to 8th nearest-neighbor sites relative to the O site ($c=1/128$). $N_{d}$ is the number of adjacent, degenerate H sites in the nearest neighbor shell, $r_{z}$ is the component of the $\mathbf{r}$ vector connecting O to H parallel to the axis of expansion of the O defect, $r$ is the magnitude of $\mathbf{r}$.
Site $N_{d}$ $_{z} rr$ $E_{f}^{\mathrm{O\textrm{-}H}}$
- (Å) (Å) (meV)
1 4 0.00 0.83 unstable
2 1 0.83 1.85 374
3a 2 1.65 2.48 118
3b 1 0.00 2.48 47.5
4 1 2.48 2.98 136
5a 2 1.65 3.41 5.3
5b 1 0.00 3.41 28.8
5c 4 3.31 3.41 75.8
6 1 0.83 3.79 -20.3
7a 2 0.00 4.14 -72.5
7b 1 3.31 4.14 34.2
7c 1 0.00 4.14 -9.0
8a 1 2.48 4.45 -2.6
8b 1 4.13 4.45 40.3
Fig. 1 presents the distribution of H interstitial sites relative to an interstitial O site and a substitutional Ti site within a $2\times 2\times 1/2$ supercell of bcc Nb. Each H site is indexed by its relative distance to the defect site. To identify the preferred H interstitial positions adjacent to an O interstitial site, we computed the formation energy of OH complexes ($E_{f}^{\mathrm{OH}}$) for all H sites up to the 8th nearest neighbor H sites in a $4\times 4\times 4$ supercell as
$E^{\mathrm{OH}}_{f}=E_{\mathrm{OH}}-E_{\mathrm{H}}-E_{\mathrm{O}}+n_{\mathrm{Nb}}\mu_{\mathrm{Nb}}\,$ (1)
where $E_{\mathrm{OH}}$, $E_{\mathrm{O}}$, and $E_{\mathrm{H}}$ are the total energies of the OH complex, a single O interstitial, and a single H insterstitial computed using density functional theory (DFT). The term $n_{Nb}$ denotes the number of Nb atoms in the supercell, and the Nb chemical potential, $\mu_{\mathrm{Nb}}$, is referenced to the total energy per atom of bcc Nb.
Table 1 summarizes the O-H complex formation energies, the degeneracies of corresponding adjacent H sites, and the position of the tetrahedral sites relative to the O interstitial. We find that the favored site is a 7th nearest neighbor tetrahedral site with a formation energy of -72.5 meV. This value is in reasonable agreement with the experimentally reported binding enthalpy of $\sim$100 meV [23]. These data show that H unambiguously occupies a well-defined tetrahedral site in the presence of an O trap. At the millikelvin operating temperatures relevant for superconducting qubits, only this site will be thermodynamically accessible, and therefore we use it when computing the tunnel splitting of O-trapped H in the remaining work.
The preference for this site can be understood on two complementary grounds. First, the elastic dipole tensors for H and O interstitials in bcc Nb take the form:
$P_{\mathrm{H}}=\begin{bmatrix}A&0&0\\ 0&B&0\\ 0&0&B\\ \end{bmatrix},P_{\mathrm{O}}=\begin{bmatrix}C&0&0\\ 0&-D&0\\ 0&0&-D\\ \end{bmatrix}$
For H, $A\approx B$ [28] implies an almost isotropic local expansion of the lattice. In contrast, O interstitials exhibit an axis of expansion and a perpendicular plane of contraction. An H atom therefore minimizes its elastic interaction energy by occupying a site within the O-induced contraction plane, consistent with the DFT results. Second, O occupies a negatively charged (O2-) octahedral site in bcc Nb. The $2p^{6}$ valence electrons of the oxide ion preferentially align with the 6 adjacent Nb sites; however, as the bcc octahedral sites are anisotropic, the Nb–O bonds distort to equalize their lengths. This results in an expansion of the bond along the short axis and compression of the bond along the long axes. The presence of H at the optimal tetrahedral sites enhances the compression of the Nb–O bond along the long axes, stabilizing the configuration relative to alternative sites. We anticipate similar mechanisms will apply to other octahedral interstitials such as carbon and nitrogen.
Table 2: Formation energy ($E_{f}$) of Ti-H, Zr-H, and Ta-H complexes in bcc Nb with H located at the 1st to 5th nearest-neighbor sites relative to the O site ($c=1/54$). $N_{d}$ is the number of adjacent, degenerate H sites in the nearest neighbor shell, $r$ is distance between the H site and the substitutional defect.
Site $N_{d}$ $r$ $E_{f}^{\mathrm{Ti\textrm{-}H}}$ $E_{f}^{\mathrm{Zr\textrm{-}H}}$ $E_{f}^{\mathrm{Ta\textrm{-}H}}$
- (Å) (meV) (meV) (meV)
1 24 1.85 -73.8 -50.0 27.1
2 2 2.98 -0.4 -39.9 10.0
3 2 3.79 4.7 -9.8 1.6
4a 6 4.45 8.4 25.3 0.5
4b 1 4.45 -0.2 -1.9 -2.83
5 4 5.01 -0.1 33.3 -15.9
@@ II Materials System and Defect Sites
Owing to their cubic symmetry, substitutional impurities such as Ti, Zr, and Ta, create significantly simpler H-site environments than interstitial O (Fig. 1). Their elastic dipole tensors are isotropic, so the long-range elastic interactions between H and a substitutional defect is governed primarily by the sign and magnitude of the defect’s elastic dipole tensor rather than strain-field anisotropy. We computed the formation energies of H bound by substitutional Ti, Zr, and Ta defects using a $3\times 3\times 3$ supercell using Eq. 1 with appropriate modifications. Table 2 summarizes the computed formation energies up to the 5th nearest neighbor sites. For both Ti-H and Zr-H complexes, the first-nearest neighbor tetrahedral site exhibits the lowest formation energy, with Ti-H having a $E_{f}$ comparable to that of the O-H complex. Zr-trapped H is slightly less stable with $E_{f}=-50.0$ meV, however its stability extends further, up to the third-nearest neighbor, in contrast to Ti-trapped H, which is essentially only stable at the first-nearest neighbor site. In contrast, Ta defects repel H. The Ta-H formation energy remains positive through the fourth-nearest neighbor sites, indicating that H is excluded within in region of radius $\sim$5 Å around Ta in Nb. Based on these results, we restrict our tunnel-splitting calculations to Ti-trapped and Zr-trapped H in the present work.
III Tunneling Model
III.1 Theory
@@
Here, we ignore the cross-coupling to the remaining phonon modes $\{\omega_{iQk}\}_{i\neq\mathbf{q},Q}$ as they weakly influence the computed tunnel splitting, as we argue below. This formulation retains the essential anharmonicity required for accurate tunneling predictions and avoids reliance on thermodynamically unstable symmetrized structures. Physically, the dominant anharmonicity arises from the hydrogen phonon modes and the composite coordinate $Q$, which parameterizes the Nb lattice distortion between degenerate configurations (Fig. 2, left) and discussed further below. The full derivation, including the transformation from the nuclear Hamiltonian and its expansion in phonon coordinates, is provided in the Appendix. Note that in the limit where $Q$ is fixed at its symmetric value, $Q_{s}$, Eq. 2 reduces to a 3D H coordinate spanning both wells–the conventional light-particle picture in a rigid lattice.
III.2 Numerical Implementation
[…]
@@ III.2 Numerical Implementation
1/16222Computed using 5D Hamiltonian. 1.32 0.069 18.4 20.7333Evaluated using grid interpolation. 275.411footnotemark: 1 245.3 - $1.9\times 10^{-3}$ -
@@
IV Results and Discussion
IV.1 Localization and Lattice Coupling
@@
The wavefunction of the first-excited state follows a similar trend; however, it undergoes a phase flip in the symmetric configuration that persists toward the right well. Fig. 2g(h) further shows that the potential energy and wavefunction along the $Q$ coordinate has a harmonic character at each H minima, $\mathbf{q}_{1}$ ($\mathbf{q_{2}}$). We also plot the image of the potential energy surface at the other H minima, $\mathbf{q}_{2}$ ($\mathbf{q}_{1}$). Both surfaces meet at common coordinate, $Q_{c}$, and potential energy, $E_{c}^{\prime}$, which defines the symmetric configuration and effective coincidence energy.
IV.2 Tunnel Splitting and Dimensionality Effects
With the full 4D Hamiltonian, we compute a tunnel splitting ($J_{\mathrm{H,D}}$) ranging from 0.31 meV ($c=1/16$) to 0.064 meV ($c=1/54$) for hydrogen (H) and 0.024 meV ($c=1/16$) to $4.8\times 10^{-3}$ meV ($c=1/54$) for deuterium (D) TLS in bcc Nb (Table 3). $J_{\mathrm{H}}$ computed at $c=1/54$ is approximately 3-fold smaller than experimental estimates derived from specific heat (0.19 meV) and neutron scattering experiments (0.21 meV) [23], consistent with later measurements reporting values ranging from 0.170 – 0.230 meV [24]. $J_{\mathrm{D}}$ computed at $c=1/54$ is about 4 times lower than the estimate from specific heat experiments (0.021 meV) [30]. Interestingly, $J_{H}$ for self-trapped hydrogen is comparable to $J_{D}$ computed for oxygen-trapped hydrogen. This arises due to the confluence of a larger H–H site separation ($R_{H}$), a larger lattice displacement ($R_{c}$) and coincidence energy ($E_{c}$) between the local and symmetric configurations, and a larger barrier ($V$) in the symmetric configuration.
@@ IV.2 Tunnel Splitting and Dimensionality Effects
Given the sensitivity of the tunnel splitting to the anharmonic terms in the Hamiltonian, we have further validated our results by comparing the computed transition frequencies between the ground state and excited states against measured inelastic neutron scattering data for O-H and H defects [19]. The reported transition frequencies are 108 (106) meV and 159 (163) meV for O-trapped H (self-trapped H), with the latter transition two-fold degenerate due to the local symmetry of tetrahedral sites. Our results reproduce these values within a few meV at both defect concentrations (Table 5). By including the composite phonon mode $Q$, we find additional transitions appear in the excitation spectrum of the 4D Hamiltonian with frequencies dictated by the nearly harmonic states of the $Q$ mode. As the energy of these excited states overlap with the excited-state energies of the H-dominant modes, assessing the appropriate transition frequencies required a manually assignment. This was performed by plotting the wavefunctions of the excited states and selecting the excited states whose nominal wavefunction was consistent with the harmonic oscillator wavefunction of the desired transition. The transition energy between the ground and first excited state varied between 103-110 meV for all evaluated defects, indicating a weak sensitivity to barrier height. The transition energy between the ground state and higher quasi-degenerate excited states ranged between 140–173 meV depending on $c$ and defect type. These results confirm that our sampled potential $V(\mathbf{q},Q)$ accurately captures the relevant O-H energy landscape and represent a clear improvement over the prior 3D Hamiltonian approach [18], likely due to the higher fidelity DFT functional choice and denser sampling.
@@
IV.3 Impact of Nb Mode Coupling
The effect of the harmonic Nb mode couplings on $J$ is expected to be small. At the symmetric configuration with a potential energy minima at $Q=Q_{s}$, the wells are degenerate, however, second-order mode couplings $\omega_{q_{i}q_{Q}}$ to the phonon bath slightly offset and lower the local minimum from $\{q_{i}\}_{\neq Q}=0$. Degenerate wells remain following this offset, which reduces the transition energy and marginally increases the tunnel splitting. We estimate the maximum reduction by comparing the energy difference between the effective, $E_{c}^{\prime}$, and true, $E_{c}$, coincidence energies when all other modes are fully relaxed at $Q=Q_{s}$. This energy difference is typically between 121-2 meV lower than the symmetric structure, corresponding to about $10-20\%$ of the total symmetrization energy. For the O-H defect, $E_{c}$ is only 1 meV lower than $E_{c}^{\prime}$; such a minor correction cannot reconcile our computed $J=0.064$ meV ($c=1/54$) with the experimental value of $J_{\mathrm{H}}=0.19$ meV. Indeed, this result follows from the fact that $E_{c}^{\prime}$ for $c=1/16$ is 5 meV lower than $E_{c}^{\prime}$ for $c=1/54$, yet $J_{\mathrm{H}}$ is only 50% larger than the experimental value.
@@ IV.3 Impact of Nb Mode Coupling
IV.4 Mass Effects
The dominant anharmonicity arises from the cross-coupling $\mathbf{q}-Q$ between the H modes and the composite Nb lattice mode. This interaction strongly influences tunneling behavior and challenges the validity of the light-particle picture. For a harmonic oscillator, the wavefunction amplitude $\psi\propto\exp{(-\omega Q^{2}/(2\hbar))}$. The mode frequency along the composite coordinate is given by $\omega=\sqrt{2E_{c}}/Q_{c}$, which leads to the ratio: $\psi(Q_{c})/\psi(0)=\exp\left(-E_{c}/\hbar\omega\right)$. By scaling $Q$ over a wide range of values, we evaluate how the mode frequency and coincidence energy affect the tunneling frequency of O-H and O-D defects. This is readily achieved by replacing $Q$ with $\sqrt{m/m_{\mathrm{Nb}}}Q$. Physically, this rescaling either enhances ($m<m_{\mathrm{Nb}}$) or suppresses ($m>m_{\mathrm{Nb}}$) tunneling by increasing (decreasing) the magnitude of the wavefunction in the symmetric configuration ($Q=Q_{s}$) and thereby increasing (decreasing) wavefunction overlap. Fig. 3 shows that tunnel splitting decreases by approximately a factor of 2-3 when moving from Nb to Ta. We also find that $J=J_{\mathrm{4D}}$ can vary by nearly an order of magnitude compared to the light-particle value ($J_{\mathrm{3D}})$, highlighting the critical role of these anharmonic couplings in renormalizing tunnel splittings. This reduction shifts the expected lower-bound resonant frequency of H-based TLS into the 1–20 GHz range. However, strain coupling energies of 0.2–1 eV imply that even small lattice strains (as low 0.002-0.02%) are sufficient to raise the resonant frequency above typical superconducting qubit operating frequencies, effectively suppressing most TLS from acting as resonant defects.
IV.5 Acoustic Phonons and Strain Interactions